# Decoupling Fine Detail and Global Geometry for Compressed Depth Map Super-Resolution

### v1

@@ Abstract
Recovering high-quality depth maps from compressed sources has gained significant attention due to the limitations of consumer-grade depth cameras and the bandwidth restrictions during data transmission. However, current methods still suffer from two challenges. First, bit-depth compression produces a uniform depth representation in regions with subtle variations, hindering the recovery of detailed information. Second, densely distributed random noise reduces the accuracy of estimating the global geometric structure of the scene. To address these challenges, we propose a novel framework, termed geometry-decoupled network (GDNet), for compressed depth map super-resolution that decouples the high-quality depth map reconstruction process by handling global and detailed geometric features separately. To be specific, we propose the fine geometry detail encoder (FGDE), which is designed to aggregate fine geometry details in high-resolution low-level image features while simultaneously enriching them with complementary information from low-resolution context-level image features. In addition, we develop the global geometry encoder (GGE) that aims at suppressing noise and extracting global geometric information effectively via constructing compact feature representation in a low-rank space. We conduct experiments on multiple benchmark datasets, demonstrating that our GDNet significantly outperforms current methods in terms of geometric consistency and detail recovery. In the ECCV 2024 AIM Compressed Depth Upsampling Challenge [2], our solution won the 1st place award. Our codes will be available.
@@
Figure 1: (a) illustrates the degradation process of obtaining a compressed depth map. To be specific, the original depth map is degraded by downsampling, bit-depth compression, and noise introduced during data acquisition and transmission. (b) presents the motivation behind the proposed GDNet. The core idea is to leverage the compressed depth map for capturing global geometric information, while utilizing the RGB image to extract detailed geometric features. As a result, our GDNet can effectively reconstruct a high-quality depth map.
@@ 1 Introduction
In summary, the main contributions of this paper are outlined as follows:
@@ 2.1 Depth map super-resolution
To overcome these limitations, convolutional neural networks (CNNs) have been widely employed to learn complex mappings from low-resolution (LR) to high-resolution (HR) depth maps [12, 30, 26]. CNN-based methods have demonstrated notable improvements in recovering fine geometry details and enhancing overall depth map quality [15, 39, 33, 34, 43]. More recently, transformer-based models have also been explored to capture long-range dependencies, leading to further advancements in DMSR [21].
@@ 2.2 Guided depth map super-resolution
Learning-based methods focus on leveraging the powerful feature extraction capabilities of convolutional neural networks (CNNs) to effectively enhance the resolution of depth maps [51, 50]. These approaches utilize high-resolution RGB images as guidance to improve the resolution of depth maps, effectively transforming low-resolution inputs into high-resolution outputs. Specifically, these GDSR models exploit the structural similarities between depth maps and RGB images, which allows for more precise edge preservation and detailed reconstruction [48]. Furthermore, advanced techniques such as attention mechanisms and feature fusion strategies have been introduced to refine the integration of RGB guidance and depth features, ultimately achieving superior performance compared to traditional methods [43].
@@ 3.1 Overall framework
Fine geometry detail encoder. To begin with, the input image is fed into the fine geometry detail encoder to extract the fine geometry details of the scene, which can be expressed as follows:
@@
Figure 3: (a) shows the detailed structure of the proposed fine geometry detail encoder (FGDE). The purpose of FGDE is to preserve fine geometric details in high-resolution low-level image features while augmenting them with supplementary information derived from low-resolution context-level image features. (b) presents the process of low-rank feature reconstruction in global geometry encoder (GGE). By integrating low-rank feature reconstruction, GGE aims at completing feature reconstruction in a low-rank space, thereby achieving the objectives of noise suppression and effective extraction of global geometric cues.
@@ 3.3 Global geometry encoder
However, we cannot directly learn a set of linearly independent vectors, which may lead to a problem: the inverse of $(B^{T}B)$ may not exist. To address this issue, we employ QR decomposition to construct a full-rank matrix, which serves as the basis vectors for the low-rank space. Specifically, we perform QR decomposition on $B$ as follows:
@@ 4.1 Experimental settings
Table 1: Quantitative performance comparison of various methods on the AIM 2024 Compressed Depth Upsampling Challenge dataset. The table reports the mean absolute error (MAE) and root mean squared error (RMSE). Our proposed method outperforms the existing approaches, achieving the lowest MAE and RMSE, indicating superior accuracy in depth map super-resolution under compression scenarios.
@@ 4.2 Quantitative results
The experimental results, as summarized in Tables 1 and 2, demonstrate the superior performance of our proposed method compared to several state-of-the-art approaches on both the AIM 2024 Compressed Depth Upsampling Challenge dataset and the synthesized Compressed-NYU dataset. On the AIM 2024 Compressed Depth Upsampling Challenge dataset, our method achieves a MAE of 0.212 and an RMSE of 0.375, significantly outperforming existing methods such as DAS-Depth (MAE: 0.294, RMSE: 0.432) and DINOv2-ControlNet (MAE: 0.498, RMSE: 0.816). Notably, our approach also delivers a remarkable improvement over simpler interpolation-based methods like Bicubic (MAE: 16.48, RMSE: 57.69). Similarly, on the synthesized Compressed-NYU dataset, our method achieves the lowest MAE of 0.0739 and RMSE of 0.0322, which is a substantial improvement compared to SGNet, the next best-performing method, with a MAE of 0.1051 and an RMSE of 0.0426. These results clearly demonstrate that our method is highly effective in depth map super-resolution under compression scenarios, showing its superiority in preserving fine geometry details and maintaining global geometric consistency.
@@ 4.3 Visual results
Table 2: Quantitative performance comparison on the synthesized Compressed-NYU dataset, evaluating different methods using mean absolute error (MAE) and root mean squared error (RMSE). The results demonstrate the effectiveness of our approach, which achieves the best performance with significantly lower error metrics compared to other well-known methods.
Method MAE $\downarrow$ RMSE $\downarrow$
Bicubic 2.6918 2.5250
Bilinear 2.6939 2.5268
Depth Anything [41] 1.2525 1.1360
Depth Anything V2-Small [42] 0.6026 0.5130
Depth Anything V2-Base [42] 0.6037 0.5237
Depth Anything V2-Large [42] 0.5994 0.5247
SGNet [38] 0.1051 0.0426
Ours 0.0739 0.0322
@@ 4.4 Ablation study
Table 3: An ablation study of our GDNet was conducted on the synthesized Compressed-NYU dataset. In this study, we assess the contributions of each module and loss function within GDNet, including FGDE, GGE, and SILog loss.
FGDE GGE L1 MSE SILog MAE $\downarrow$ RMSE $\downarrow$
0.1079 0.0488
0.4949 0.3781
0.0796 0.0347
0.0796 0.0372
0.0739 0.0322

### v4

@@ Abstract
Recovering high-quality depth maps from compressed sources has gained significant attention due to the limitations of consumer-grade depth cameras and the bandwidth restrictions during data transmission. However, current methods still suffer from two challenges. First, bit-depth compression produces a uniform depth representation in regions with subtle variations, hindering the recovery of detailed information. Second, densely distributed random noise reduces the accuracy of estimating the global geometric structure of the scene. To address these challenges, we propose a novel framework, termed geometry-decoupled network (GDNet), for compressed depth map super-resolution that decouples the high-quality depth map reconstruction process by handling global and detailed geometric features separately. To be specific, we propose the fine geometry detail encoder (FGDE), which is designed to aggregate fine geometry details in high-resolution low-level image features while simultaneously enriching them with complementary information from low-resolution context-level image features. In addition, we develop the global geometry encoder (GGE) that aims at suppressing noise and extracting global geometric information effectively via constructing compact feature representation in a low-rank space. We conduct experiments on multiple benchmark datasets, demonstrating that our GDNet significantly outperforms current methods in terms of geometric consistency and detail recovery. In the ECCV 2024 AIM Compressed Depth Upsampling Challenge, our solution won the 1st place award. Our codes are available at: https://github.com/Ian0926/GDNet.
@@
Figure 1: (a) Illustration of the Degradations during Data Acquisition and Transmission. To be specific, the quality of the depth map is compromised due to downsampling, bit-depth compression, and noise introduced during data acquisition and transmission. (b) The Motivation behind the Proposed GDNet. The core idea is to leverage the compressed depth map for capturing global geometric information, while utilizing the RGB image to extract detailed geometric features. As a result, our GDNet can effectively reconstruct a high-quality depth map.
@@ 1 Introduction
The contributions of this paper are outlined as follows:
@@ 2.1 Depth map super-resolution
To overcome these limitations, convolutional neural networks (CNNs) have been widely employed to learn complex mappings from low-resolution (LR) to high-resolution (HR) depth maps [13, 32]. CNN-based methods have demonstrated notable improvements in recovering fine geometry details and enhancing overall depth map quality [16, 40, 34, 45, 38]. More recently, transformer-based models have also been explored to capture long-range dependencies, leading to further advancements in DMSR [22].
@@ 2.2 Guided depth map super-resolution
Learning-based methods focus on leveraging the powerful feature extraction capabilities of convolutional neural networks (CNNs) to effectively enhance the resolution of depth maps [52]. These approaches utilize high-resolution RGB images as guidance to improve the resolution of depth maps, effectively transforming low-resolution inputs into high-resolution outputs. Specifically, these GDSR models exploit the structural similarities between depth maps and RGB images, which allows for more precise edge preservation and detailed reconstruction [50]. Furthermore, advanced techniques such as attention mechanisms and feature fusion strategies have been introduced to refine the integration of RGB guidance and depth features, ultimately achieving superior performance [45].
@@ 3.1 Overall framework
Fine geometry detail encoder. To begin with, the input image is fed into the fine geometry detail encoder to extract the fine geometry details, which can be expressed as follows:
@@ 3.2 Fine geometry detail encoder
Figure 3: (a) The Detailed Structure of the Proposed Fine Geometry Detail Encoder (FGDE). The purpose of FGDE is to preserve fine geometric details in high-resolution low-level image features while augmenting them with supplementary information derived from low-resolution context-level image features. (b) The Process of Low-rank Feature Reconstruction in Global Geometry Encoder (GGE). By integrating low-rank feature reconstruction, GGE aims at completing feature reconstruction in a low-rank space, thereby achieving the objectives of noise suppression and effective extraction of global geometric cues.
@@ 3.3 Global geometry encoder
However, it is difficult to directly learn a set of linearly independent vectors, which may lead to a problem: the inverse of $(B^{T}B)$ may not exist. To address this issue, we employ QR decomposition to construct a full-rank matrix, which serves as the basis vectors for the low-rank space. Specifically, we perform QR decomposition on $B$ as follows:
@@ 4.1 Experimental settings
Table 1: Quantitative Performance Comparison of Various Methods on the AIM 2024 Compressed Depth Upsampling Challenge Dataset. The table reports both MAE and RMSE. Our proposed method outperforms the existing approaches, achieving the lowest MAE and RMSE, indicating superior accuracy in depth map super-resolution under compression scenarios.
@@
Table 2: Quantitative Performance Comparison on the Synthesized Compressed-NYU Dataset, evaluating different methods using both MAE and RMSE. The results demonstrate the effectiveness of our approach, which achieves the best performance with significantly lower error metrics compared to other methods.
Method MAE $\downarrow$ RMSE $\downarrow$
Depth Anything [43] 1.1360 1.2525
Depth Anything V2-Small [44] 0.5130 0.6026
Depth Anything V2-Base [44] 0.5237 0.6037
Depth Anything V2-Large [44] 0.5247 0.5994
UniDepth V1 [28] 0.1702 0.2404
UniDepth V2 [28] 0.1414 0.1944
iDisc [27] 0.2735 0.3778
AdaBins [2] 0.4233 0.5694
SGNet [39] 0.0426 0.1051
Ours 0.0322 0.0739
@@ 4.2 Quantitative results
The experimental results, as summarized in Tables 1 and 2, demonstrate the superior performance of our proposed method compared to several state-of-the-art approaches on both the AIM 2024 Compressed Depth Upsampling Challenge dataset and the synthesized Compressed-NYU dataset. On the AIM 2024 Compressed Depth Upsampling Challenge dataset, our method achieves a MAE of 0.212 and an RMSE of 0.375, significantly outperforming existing methods such as DAS-Depth (MAE: 0.294, RMSE: 0.432) and DINOv2-ControlNet (MAE: 0.498, RMSE: 0.816). Notably, our approach also delivers a remarkable improvement over simpler interpolation-based methods like Bicubic (MAE: 16.48, RMSE: 57.69). Similarly, on the synthesized Compressed-NYU dataset, our method achieves the lowest MAE of 0.0322 and RMSE of 0.739, which is a substantial improvement compared to SGNet, the next best-performing method, with a MAE of 0.0426 and an RMSE of 0.1051. These results clearly demonstrate that our method is highly effective in depth map super-resolution under compression scenarios, showing its superiority in preserving fine geometry details and maintaining global geometric consistency.
@@ 4.3 Visual results
Table 3: An Ablation Study of Our GDNet Conducted on the Synthesized Compressed-NYU Dataset. In this study, we assess the contributions of each module and loss function within GDNet.
FGDE DCPM GGE LFR L1 MSE SILog MAE $\downarrow$ RMSE $\downarrow$
0.0488 0.1079
0.0475 0.0974
0.3781 0.4949
0.0566 0.1048
0.0347 0.0796
0.0372 0.0796
0.0322 0.0739
@@ 4.4 Ablation study
Table 4: An Ablation Study of the Number of CA and SA. $N_{SA}$ and $N_{CA}$ denotes the numbers of SA and CA, respectively.
Method $N_{SA}$ $N_{CA}$ MAE $\downarrow$ RMSE $\downarrow$
$\text{S}_{2}\text{C}_{1}$ 2 1 0.0357 0.0789
$\text{S}_{3}\text{C}_{1}$ 3 1 0.0385 0.0829
$\text{S}_{1}\text{C}_{2}$ 1 2 0.0345 0.0765
$\text{S}_{1}\text{C}_{3}$ 1 3 0.0372 0.0796
Ours 1 1 0.0322 0.0739
Number of CA and SA. We have analyzed the impact of varying the number of SA and CA on performance in Table 4. We select the best setting for our GDNet.